OpenStudy (anonymous):

Evaluate the integral: e^(5x)cos(e^5x) with an upper limit of 3 and lower limit of 0

OpenStudy (dumbcow):

substitution: u = e^5x du = 5*e^5x dx --> dx = du/5e^5x --> integral (1/5) cos(u) du

OpenStudy (anonymous):

I got sin e^5-sin 1/5 but I think it is supposed to be a decimal number i have no idea how to go about this

OpenStudy (dumbcow):

i think it should be sin(e^15) ....5x evaluated at x=3 yes i get 0.029, just use a calculator to get approximate answer, should be in radians

OpenStudy (anonymous):

got it, thanks. I see where I went wrong.